A (yearly) cash flow stream is X = (-65, 21, 21, 21, 21, 21, 21). The spot rates (yearly in percentage) are s = (5.0, 5.3, 5.6, 5.8, 6.0, 6.1).

(a) Find the current discount factors dk.

(b) Use the discount factors to determine the (net) present value of the cash flow stream. (Keep your answer to 2 decimal places, e.g. xxx.12)
Answer #1
Year Cash flow Spot rate a...Discount Factors PV of C/f
1 2 3 4 5 6=2*5
0 -65 1 -65
1 21 0.05 1/1.05^1= 0.95238 20
2 21 0.053 1/1.053^2= 0.90187 18.9392
3 21 0.056 1/1.056^3= 0.84920 17.8331
4 21 0.058 1/1.058^4= 0.79810 16.7601
5 21 0.06 1/1.06^5 = 0.74726 15.6924
6 21 0.061 1/1/061^6= 0.70098 14.7207
b..Net Present value of the C/f stream = 38.9455
ie. 38.95
ANSWERS: Discount factors:
1
0.95238
0.90187
0.84920
0.79810
0.74726
0.70098
b..Net Present value of the C/f stream = 38.95